# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ Appendix A Proofs
$\xi_{1,j}$ $=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(\Sigma_{t}^{1/2}Q_{i})(\Sigma_{t}^{1/2}Q_{i})^{\top}\Sigma_{t}^{-1}\check{a}$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(e_{j}^{\top}\Sigma_{t}^{1/2}Q_{i})(Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a})$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a})(e_{j}^{\top}\Sigma_{t}^{1/2}Q_{i})$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i},\quad A_{j}:=\Sigma_{t}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\,,$
@@
$\|\check{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq\|\check{a}-\xi_{1}\|_{\infty}+\|\xi_{2}\|_{\infty}=\max_{1\leq j\leq p+q}\left|\check{a}_{j}-\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i}\right|+\|\xi_{2}\|_{\infty}\,.$ (A.7)
@@
$\left\|A_{j}\right\|_{F}^{2}=\operatorname{tr}\left(A_{j}^{\top}A_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{t}^{-\frac{1}{2}}\Sigma_{t}^{-\frac{1}{2}}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(\Sigma_{t}^{\frac{1}{2}}e_{j}\check{a}^{\top}\Sigma_{t}^{-1}\check{a}e_{j}^{\top}\Sigma_{t}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(e_{j}\Sigma_{t}^{\frac{1}{2}}\Sigma_{t}^{\frac{1}{2}}\check{a}^{\top}\Sigma_{t}^{-1}\check{a}\right)$
$=\left(\Sigma_{t}\right)_{jj}\check{a}^{\top}\Sigma^{-1}_{t}\check{a}\leqslant V_{1}^{\prime}S_{1}\,,$
where the bound in last line follows form Assumption 5.6 and Lemma A.1. Furthermore, observe that conditionally on treatment indicators and the observations in $\mathcal{D}_{1}$ (call this sigma field $\mathcal{F}_{0,n}$), $Q_{i}^{\top}A_{j}Q_{i}$ (for $i$ with $A_{i}=1$) are i.i.d with mean $\check{a}_{j}$. Hence, an application of Corollary 2.8 of Zajkowski, (2020) (see Lemma A.2 for details) yields:
$\mathbb{E}\left[{\sf exp}\left\{t\left(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j}\right)\right\}\mid\mathcal{F}_{0,n}\right]\leqslant{\sf exp}\left\{8C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\right\},\quad\forall t\leqslant\frac{C_{2}}{2\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}}.$ (A.8)
@@
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j}\right\}\mid\mathcal{F}_{0,n}\right]$ $=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
=$=\prod_{\{i:A_{i}=1\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}(Q_{i}^{\top}A_{j}Q_{i}-\check{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$ $\leq\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}\left(\xi_{1}-\check{a}\right)_{j}\geqslant 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j}\geqslant 9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left({\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j})}\geq{\sf exp}{(9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mid\mathcal{F}_{0,n}\right)$
$\leq\sum_{j=1}^{p+q}{\sf exp}{(-9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\check{a}\right)_{j})}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$ $=\mathbb{P}\left(\|\xi_{1}-\check{a}\|_{\infty}\geq 9C_{1}\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)$
@@
Defining $9C_{1}\iota^{2}\sqrt{2V_{1}S_{1}}=K_{4}$ and taking expectation with respect to the $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (A.2), we conclude that $\|\xi_{1}-\check{a}\|_{\infty}\leq K_{4}\sqrt{\log{(p+q)}/n_{t}}$ with probability approaching to 1.
@@
$\|\xi_{2}\|_{\infty}\leq\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(\mathbf{W}_{t,i}-\mu_{t})^{\top}\Sigma_{t}^{-1}\check{a}\right|=\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a}\right|:=\left\|\mu_{t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Z_{t,i}\right|\,,$
where $Z_{t,i}=Q_{i}^{\top}\Sigma_{t}^{-1/2}\check{a}$. Note that $Z_{t,i}\in\mathbb{R}$ and therefore, by Chebychev’s inequality, we have
@@
Now $\operatorname{Var}(Z_{t}\mid A=1)=\check{a}^{\top}\Sigma_{t}^{-1}\check{a}$ which is upper bounded by $V_{1}^{\prime}$ by Lemma A.1. Therefore, taking $t=\sqrt{\log{(p+q)}/n_{t}}$ we obtain
@@
Furthermore, $\|\mu_{t}\|_{\infty}\leq C_{t}$ by Assumption 5.6. Hence, we conclude that $\|\xi_{2}\|_{\infty}\leq K_{5}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.3) and (A.7) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{1}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1, for some constant $K_{1}>0$.
Next, we show the second part of the feasibility of the optimization problem (5.4), that is $\|\tau_{1}^{*}\|_{\infty}\leq n_{t}^{-2/3}$. Let $z=\Sigma_{t}^{-1/2}\check{a}$, then $\|z\|_{2}\leq V_{1}^{{}^{\prime}1/2}$ by Lemma A.1. Let $\mathbf{z}=z/\|z\|_{2}$, and as $Q_{i}$’s are sub-Gaussian and by our design (Assumption 5.6), we can write:
[…]
@@
[…]
@@
We now demonstrate that $\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu^{\top}\Sigma_{c}^{-1}(\mathbf{X}_{c,i}-\mu_{c})$ for $1\leq i\leq n$ is a feasible solution to the optimization problem (5.5) with high probability, i.e.,
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leqslant K_{5}\sqrt{\log(p)/n_{c}}\,,\ \ \|\tau_{2}^{*}\|_{\infty}\leq n_{c}^{-2/3}\,.$ (A.12)
@@
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\bar{X}-\mu\|_{\infty}+\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\,.$ (A.13)
For the first term, we have $\|\bar{X}-\mu\|_{\infty}=K_{6}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1 for some constant $K_{6}>0$, by a simple application of Hoeffding’s inequality along with a union bound. We next obtain the upper bound of $\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}$ in Equation (A.13). Write $(\mathbf{X}_{c}^{\top}\tau_{2}^{*})$ as:
$\mathbf{X}_{c}^{\top}\tau_{2}^{*}$ =$=\sum_{\{i:A_{i}=0\}}X_{c,i}\tau_{2,i}^{*}=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}X_{c,i}^{\top}(X_{c,i}-\mu_{c})\Sigma_{c}^{-1}\mu$
$=\underbrace{\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu}_{:=\xi_{3}}+\underbrace{\mu_{c}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu}_{:=\xi_{4}}\,.$ (A.14)
@@
$\xi_{3,j}$ $=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{c})(X_{c,i}-\mu_{c})^{\top}\Sigma_{c}^{-1}\mu$
$=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(\Sigma_{c}^{1/2}P_{i})(\Sigma_{c}^{1/2}P_{i})^{\top}\Sigma_{c}^{-1}\mu$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(e_{j}^{\top}\Sigma_{c}^{1/2}P_{i})(P_{i}^{\top}\Sigma_{c}^{-1/2}\mu)$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}\Sigma_{c}^{-1/2}\mu)(e_{j}^{\top}\Sigma_{c}^{1/2}P_{i})$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}G_{j}P_{i},\quad G_{j}=\Sigma_{c}^{-\frac{1}{2}}\mu e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\,.$
@@
$\|\mu-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\mu-\xi_{3}\|_{\infty}+\|\xi_{4}\|_{\infty}=\max_{1\leq j\leq p}\left|\mu_{j}-\frac{1}{n_{c}}\sum_{i:A_{i}=0}P_{i}^{\top}G_{j}P_{i}\right|+\|\xi_{4}\|_{\infty}\,.$ (A.15)
@@
$\|G_{j}\|_{F}^{2}=tr\left(G_{j}^{\top}G_{j}\right)$ $=tr\left(\Sigma_{c}^{\frac{1}{2}}e_{j}\mu^{\top}\Sigma_{c}^{-\frac{1}{2}}\Sigma_{c}^{-\frac{1}{2}}\mu e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\right)$
$=tr\left(e_{j}^{\top}\Sigma_{c}^{\frac{1}{2}}\Sigma_{c}^{\frac{1}{2}}e_{j}\mu^{\top}\Sigma_{c}^{-1}\mu\right)$
$=(\Sigma_{c})_{jj}\mu^{\top}\Sigma_{c}^{-1}\mu\leqslant V_{2}S_{2}\,,$
where the bound in last line follows Assumption 5.6 and Lemma A.1. Note that $P_{i}^{\top}G_{j}P_{i}$ (for $i$ with $A_{i}=0$) are i.i.d with mean $\mu$. Thus, an application of Hanson-Wright inequality (Theorem 1.1 of Rudelson and Vershynin, (2013)) leads to:
$\mathbb{E}\left[{\sf exp}\left\{t\left(P_{i}^{\top}G_{j}P_{i}-\mu_{j}\right)\right\}\mid A_{i}=0\right]\leqslant{\sf exp}\left\{C_{3}t^{2}\varsigma^{4}V_{2}S_{2}\right\},\quad\forall t\leqslant\frac{C_{4}}{\varsigma^{2}\sqrt{V_{2}S_{2}}}.$ (A.16)
@@
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{c}}t\left(\xi_{3}-\mu\right)_{j}\right\}\mid A_{1:n}\right]$ $=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}G_{j}P_{j}-\mu_{j})\right\}\mid A_{1:n}\right]$
=$=\prod_{\{i:A_{i}=0\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}(P_{i}^{\top}G_{j}P_{j}-\mu_{j})\right\}\mid A_{j}=0\right]$

### v3

@@ Appendix A Proofs
$\xi_{1,j}$ $=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(W_{t,i}-\mu_{W,t})(W_{t,i}-\mu_{W,t})^{\top}\Sigma_{W,t}^{-1}\tilde{a}$
$=\frac{1}{n_{t}}e_{j}^{\top}\sum_{\{i:A_{i}=1\}}(\Sigma_{W,t}^{1/2}Q_{i})(\Sigma_{W,t}^{1/2}Q_{i})^{\top}\Sigma_{W,t}^{-1}\tilde{a}$
[…]
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}\Sigma_{W,t}^{-1/2}\tilde{a})(e_{j}^{\top}\Sigma_{W,t}^{1/2}Q_{i})$
$=\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i},\quad A_{j}:=\Sigma_{W,t}^{-\frac{1}{2}}\tilde{a}e_{j}^{\top}\Sigma_{W,t}^{\frac{1}{2}}\,$,
@@
$\|\tilde{a}-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq\|\tilde{a}-\xi_{1}\|_{\infty}+\|\xi_{2}\|_{\infty}=\max_{1\leq j\leq p+q}\left|\tilde{a}_{j}-\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}A_{j}Q_{i}\right|+\|\xi_{2}\|_{\infty}\,.$ (A.6)
@@
$\left\|A_{j}\right\|_{F}^{2}=\operatorname{tr}\left(A_{j}^{\top}A_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{W,t}^{\frac{1}{2}}e_{j}\tilde{a}^{\top}\Sigma_{W,t}^{-\frac{1}{2}}\Sigma_{W,t}^{-\frac{1}{2}}\tilde{a}e_{j}^{\top}\Sigma_{W,t}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(\Sigma_{W,t}^{\frac{1}{2}}e_{j}\tilde{a}^{\top}\Sigma_{W,t}^{-1}\tilde{a}e_{j}^{\top}\Sigma_{W,t}^{\frac{1}{2}}\right)$
[…]
$=\left(\Sigma_{W,t}\right)_{jj}\tilde{a}^{\top}\Sigma^{-1}_{W,t}\tilde{a}\leqslant V_{1}^{\prime}S_{1}\,$,
where the bound in last line follows form Assumption 5.7, 5.8 and Lemma A.1. Furthermore, observe that conditionally on treatment indicators and the observations in $\mathcal{D}_{1}$ (call this sigma field $\mathcal{F}_{0,n}$), $Q_{i}^{\top}A_{j}Q_{i}$ (for $i$ with $A_{i}=1$) are i.i.d with mean $\tilde{a}_{j}$. Hence, an application of Corollary 2.8 of Zajkowski, (2020) (see Lemma A.2 for details) yields:
$\mathbb{E}\left[{\sf exp}\left\{t\left(Q_{i}^{\top}A_{j}Q_{i}-\tilde{a}_{j}\right)\right\}\mid\mathcal{F}_{0,n}\right]\leqslant{\sf exp}\left\{8C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\right\},\quad\forall t\leqslant\frac{C_{2}}{2\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}}.$ (A.7)
@@
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{t}}t\left(\xi_{1}-\tilde{a}\right)_{j}\right\}\mid\mathcal{F}_{0,n}\right]$ $=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}\sum_{\{i:A_{i}=1\}}(Q_{i}^{\top}A_{j}Q_{i}-\tilde{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
$=\prod_{\{i:A_{i}=1\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{t}}}(Q_{i}^{\top}A_{j}Q_{i}-\tilde{a}_{j})\right\}\mid\mathcal{F}_{0,n}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\tilde{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$ $\leq\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}\left(\xi_{1}-\tilde{a}\right)_{j}\geqslant 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left(\sqrt{n_{t}}t\left(\xi_{1}-\tilde{a}\right)_{j}\geqslant 9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\sum_{j=1}^{p+q}\mathbb{P}\left({\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\tilde{a}\right)_{j})}\geq{\sf exp}{(9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mid\mathcal{F}_{0,n}\right)$
$\leq\sum_{j=1}^{p+q}{\sf exp}{(-9C_{1}^{2}t^{2}\iota^{4}V_{1}^{\prime}S_{1})}\mathbb{E}\left[{\sf exp}{(\sqrt{n_{t}}t\left(\xi_{1}-\tilde{a}\right)_{j})}\right]$
@@
$\mathbb{P}\left(\sqrt{n_{t}}\|\xi_{1}-\tilde{a}\|_{\infty}\geq 9C_{1}^{2}t\iota^{4}V_{1}^{\prime}S_{1}\mid\mathcal{F}_{0,n}\right)$
$=\mathbb{P}\left(\|\xi_{1}-\tilde{a}\|_{\infty}\geq 9C_{1}\iota^{2}\sqrt{2V_{1}^{\prime}S_{1}}\sqrt{\frac{\log{(p+q)}}{n_{t}}}\mid\mathcal{F}_{0,n}\right)$
@@
Defining $9C_{1}\iota^{2}\sqrt{2V_{1}S_{1}}=K_{4}$ and taking expectation with respect to the $\sigma$-field $\mathcal{F}_{0,n}$ on the both side of Equation (A.2), we conclude that $\|\xi_{1}-\tilde{a}\|_{\infty}\leq K_{4}\sqrt{\log{(p+q)}/n_{t}}$ with probability approaching to 1.
@@
[…]
$=\left\|\mu_{W,t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Q_{i}^{\top}\Sigma_{W,t}^{-1/2}\tilde{a}\right|:=\left\|\mu_{W,t}\right\|_{\infty}\left|\frac{1}{n_{t}}\sum_{\{i:A_{i}=1\}}Z_{t,i}\right|\,$,
where $Z_{t,i}=Q_{i}^{\top}\Sigma_{W,t}^{-1/2}\tilde{a}$. Note that $Z_{t,i}\in\mathbb{R}$ and therefore, by Chebychev’s inequality, we have
@@
Now $\operatorname{Var}(Z_{t}\mid A=1)=\tilde{a}^{\top}\Sigma_{W,t}^{-1}\tilde{a}$ which is upper bounded by $V_{1}^{\prime}$ by Lemma A.1. Therefore, taking $t=\sqrt{\log{(p+q)}/n_{t}}$ we obtain
@@
Furthermore, $\|\mu_{W,t}\|_{\infty}\leq C_{t}$ by Assumption 5.8. Hence, we conclude that $\|\xi_{2}\|_{\infty}\leq K_{5}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1. Finally, combining this bound along with the upper bounds in Equation (A.3) and (A.6) we conclude that $\|a-\mathbf{W}_{t}^{\top}\tau_{1}^{*}\|_{\infty}\leq K_{1}\sqrt{\log{(p+q)}/n_{t}}$ with probability going to 1, for some constant $K_{1}>0$.
Next, we show the second part of the feasibility of the optimization problem (5.4), that is $\|\tau_{1}^{*}\|_{\infty}\leq n_{t}^{-2/3}$. Let $z=\Sigma_{W,t}^{-1/2}\tilde{a}$, then $\|z\|_{2}\leq V_{1}^{{}^{\prime}1/2}$ by Lemma A.1. Let $\mathbf{z}=z/\|z\|_{2}$, and as $Q_{i}$’s are sub-Gaussian and by our design (Assumption 5.7), we can write:
$\|\tau_{1}^{*}\|_{\infty}=\frac{1}{n_{t}}\|\tilde{a}^{\top}\Sigma_{W,t}^{-1}(W_{t,i}-\mu_{W,t})\|_{\infty}$ $=\frac{1}{n_{t}}\underset{i}{\max}|\tilde{a}^{\top}\Sigma_{W,t}^{-1/2}Q_{i}|$
$=\frac{1}{n_{t}}\|z\|_{2}\underset{i}{\max}|\mathbf{z}^{\top}Q_{i}|$
$\leq\frac{1}{n_{t}}V_{1}^{{}^{\prime}1/2}\underset{i}{\max}|\mathbf{z}^{\top}Q_{i}|\,.$
@@
$\mathbb{P}\left(\frac{1}{n_{t}}V_{1}^{{}^{\prime}1/2}\underset{i}{\max}|\mathbf{z}^{\top}Q_{i}|\geq t\mid\mathcal{F}_{0,n}\right)$ $=\mathbb{P}\left(\underset{i}{\max}|\mathbf{z}^{\top}Q_{i}|\geq V_{1}^{{}^{\prime}-1/2}tn_{t}\mid\mathcal{F}_{0,n}\right)$
$\leq\underset{i}{\sum}\mathbb{P}\left(|\mathbf{z}^{\top}Q_{i}|\geq V^{{}^{\prime}-1/2}n_{t}^{1/3}\mid\mathcal{F}_{0,n}\right)\hskip 7.22743pt\left[t\leftarrow n_{t}^{-\frac{2}{3}}\right]$
@@
We now demonstrate that $\tau_{2,i}^{*}=\frac{1}{n_{c}}\mu_{X}^{\top}\Sigma_{X,c}^{-1}(X_{c,i}-\mu_{X,c})$ for $1\leq i\leq n$ is a feasible solution to the optimization problem (5.5) with high probability, i.e.,
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leqslant K_{2}\sqrt{\log(p)/n_{c}}\,,\ \ \|\tau_{2}^{*}\|_{\infty}\leq n_{c}^{-2/3}\,.$ (A.11)
@@
$\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\bar{X}-\mu_{X}\|_{\infty}+\|\mu_{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\,.$ (A.12)
For the first term, we have $\|\bar{X}-\mu_{X}\|_{\infty}=K_{6}\sqrt{\log{(p)}/n_{c}}$ with probability going to 1 for some constant $K_{6}>0$, by a simple application of Hoeffding’s inequality along with a union bound. We next obtain the upper bound of $\|\mu_{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}$ in Equation (A.12). Write $(\mathbf{X}_{c}^{\top}\tau_{2}^{*})$ as:
$\mathbf{X}_{c}^{\top}\tau_{2}^{*}$
$=\sum_{\{i:A_{i}=0\}}X_{c,i}\tau_{2,i}^{*}=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}X_{c,i}^{\top}(X_{c,i}-\mu_{X,c})\Sigma_{X,c}^{-1}\mu_{X}$
$=\underbrace{\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{X,c})(X_{c,i}-\mu_{X,c})^{\top}\Sigma_{X,c}^{-1}\mu_{X}}_{:=\xi_{3}}+\underbrace{\mu_{X,c}\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{X,c})^{\top}\Sigma_{X,c}^{-1}\mu_{X}}_{:=\xi_{4}}\,.$ (A.13)
@@
$\xi_{3,j}$ $=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(X_{c,i}-\mu_{X,c})(X_{c,i}-\mu_{X,c})^{\top}\Sigma_{X,c}^{-1}\mu_{X}$
$=\frac{1}{n_{c}}e_{j}^{\top}\sum_{\{i:A_{i}=0\}}(\Sigma_{X,c}^{1/2}P_{i})(\Sigma_{X,c}^{1/2}P_{i})^{\top}\Sigma_{X,c}^{-1}\mu_{X}$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(e_{j}^{\top}\Sigma_{X,c}^{1/2}P_{i})(P_{i}^{\top}\Sigma_{X,c}^{-1/2}\mu_{X})$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}\Sigma_{X,c}^{-1/2}\mu_{X})(e_{j}^{\top}\Sigma_{X,c}^{1/2}P_{i})$
$=\frac{1}{n_{c}}\sum_{\{i:A_{i}=0\}}P_{i}^{\top}G_{j}P_{i},\quad G_{j}=\Sigma_{X,c}^{-\frac{1}{2}}\mu_{X}e_{j}^{\top}\Sigma_{X,c}^{\frac{1}{2}}\,.$
@@
$\|\mu_{X}-\mathbf{X}_{c}^{\top}\tau_{2}^{*}\|_{\infty}\leq\|\mu_{X}-\xi_{3}\|_{\infty}+\|\xi_{4}\|_{\infty}=\max_{1\leq j\leq p}\left|\mu_{X,j}-\frac{1}{n_{c}}\sum_{i:A_{i}=0}P_{i}^{\top}G_{j}P_{i}\right|+\|\xi_{4}\|_{\infty}\,.$ (A.14)
@@
$\|G_{j}\|_{F}^{2}=\operatorname{tr}\left(G_{j}^{\top}G_{j}\right)$ $=\operatorname{tr}\left(\Sigma_{X,c}^{\frac{1}{2}}e_{j}\mu_{X}^{\top}\Sigma_{X,c}^{-\frac{1}{2}}\Sigma_{X,c}^{-\frac{1}{2}}\mu_{X}e_{j}^{\top}\Sigma_{X,c}^{\frac{1}{2}}\right)$
$=\operatorname{tr}\left(e_{j}^{\top}\Sigma_{X,c}^{\frac{1}{2}}\Sigma_{X,c}^{\frac{1}{2}}e_{j}\mu_{X}^{\top}\Sigma_{X,c}^{-1}\mu_{X}\right)$
$=(\Sigma_{X,c})_{jj}\mu_{X}^{\top}\Sigma_{X,c}^{-1}\mu_{X}\leqslant V_{2}S_{2}\,$,
where the bound in last line follows Assumption 5.7, 5.8 and Lemma A.1. Note that $P_{i}^{\top}G_{j}P_{i}$ (for $i$ with $A_{i}=0$) are i.i.d with mean $\mu_{X}$. Thus, an application of Hanson-Wright inequality (Theorem 1.1 of Rudelson and Vershynin, (2013)) leads to:
$\mathbb{E}\left[{\sf exp}\left\{t\left(P_{i}^{\top}G_{j}P_{i}-\mu_{X,j}\right)\right\}\mid A_{i}=0\right]\leqslant{\sf exp}\left\{C_{3}t^{2}\varsigma^{4}V_{2}S_{2}\right\},\quad\forall t\leqslant\frac{C_{4}}{\varsigma^{2}\sqrt{V_{2}S_{2}}}.$ (A.15)
@@
$\mathbb{E}\left[{\sf exp}\left\{\sqrt{n_{c}}t\left(\xi_{3}-\mu_{X}\right)_{j}\right\}\mid A_{1:n}\right]$
$=\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}\sum_{\{i:A_{i}=0\}}(P_{i}^{\top}G_{j}P_{j}-\mu_{X,j})\right\}\mid A_{1:n}\right]$
$=\prod_{\{i:A_{i}=0\}}\mathbb{E}\left[{\sf exp}\left\{\frac{t}{\sqrt{n_{c}}}(P_{i}^{\top}G_{j}P_{j}-\mu_{X,j})\right\}\mid A_{j}=0\right]$